calculate the force of gravity of the husband and wife, the husband has a mass of 85kg and the wife has a mass of 65kg. they are 0.75 m apart



1
Expert's answer
2021-12-14T13:33:22-0500

"F=G\frac{Mm}{r^2}=6.67\cdot10^{-11}\cdot\frac{85\cdot65}{0.75^2}=6.6\cdot10^{-7}\ (N)"